OpenStudy (anonymous):

Mass number has no units, but when we use the eqn n= m/M, we say that m is in g. Why is that?

OpenStudy (anonymous):

My text book says that it is expressed in grams, but doesn't say whether it actually has units.... I'm so confused

OpenStudy (frostbite):

First of all the mass number isn't the "true" mass of an atom (way better to call the mass number for nucleon number). Second is that there is a relationship between the mass of an atom/molecule when we use the atomic mass unit \(\mu\) which we define as: \[\large 1 ~ g =1 \mu \times N _{A}\] So you see we get a relationship between our mass unit and the amount of molecules.

OpenStudy (frostbite):

The nucleon number / mass number don't have any unit that is right, but it is not about how much something weight (why I hate the name "mass number). It is about how many particles there are in the atomic core. See the difference? :)

OpenStudy (anonymous):

Hmm ok, so the relative mass number is like an average of the number of neutrons in the isotopes of that element, relative to The number of something to do with carbon 12?

OpenStudy (frostbite):

You are mixing stuff right now. The relative mass of an element is given by multiplying the different isotopes mass with their relative frequency of appearance.

OpenStudy (frostbite):

Just don't want you to get more confused :)

OpenStudy (anonymous):

Yes, I remember that now. So, how do we get a mass number for an element again?

OpenStudy (frostbite):

Can we make an agreement?

OpenStudy (anonymous):

I think so haha

OpenStudy (frostbite):

you NEVER ever say mass number, but nucleon number :P

OpenStudy (anonymous):

I have never heard of nucleon number before!

OpenStudy (anonymous):

So mass number is not to do with mass but a number of particles?

OpenStudy (anonymous):

It still doesn't make sense that we use the nucleon number in the eqn n=m/M in terms of grams!

OpenStudy (anonymous):

If the nucleon number has no units

OpenStudy (frostbite):

haha look it all become more understand: Lets look at heliums 2 natural isotopes.\[\LARGE _{2}^{3}He\] and \[\LARGE _{2}^{4}He\] The nucleon number is the number of particles so for 3-helium we can draw it as:  You can see the nucleon number (the number of particles in the NUCLEUS (smart huh :P)) is 3, BUT here come the interesting thing: The protons and neutrons don't weight the same and we have not even taken the electrons into consideration. So the MASS of the isotopes is: \[\large m(He)=2 \times m(proton) + 1 \times m(neutron) + 2 \times m(electron)\]